OpenStudy (anonymous):

Find critical values at g(θ) = 36θ − 9 tan θ

OpenStudy (dumbcow):

theta = pi/3, 2pi/3, 4pi/3, 5pi/3

OpenStudy (anonymous):

differentiating g'theta = 36 - 9 sec^2 theta sec^2 theta = 4 sec theta = +- 2 cos theta = + - 1/2 theta = pi/3 , 2pi/3, 4pi/3 , 5pi/3 as dumbcrow said